# Deblur-Avatar: Animatable Avatars from Motion-Blurred Monocular Videos

### v1

@@ 4.2 Baseline Comparisons
The baselines include two types: 1) state-of-the-art human avatar reconstruction methods, e.g. NeRF-based methods HumanNeRF [85] and ARAH [82], 3DGS-based methods GauHuman [16], GoMAvatar [84], 3DGS-Avatar [62] and GART [35] and 2) an image-space baseline that uses pre-trained video deblurring [56] for preprocessing and then trains the human avatar baselines with the deblurred inputs. These baselines are compared under the monocular setup on ZJU-MoCap-Blur and Real-Human-Blur. All experiments are conducted on NVIDIA RTX 3090 GPU.
4.3 Qualitative Results
@@
In Fig. 3, we compare our method with existing human avatar modeling methods on ZJU-MoCap-Blur. In Fig. 4, we show the novel pose synthesis results on Real-Human-Blur. One can observe that current human avatar modeling methods can not recover sharp details from motion blur and our method Deblur-Avatar outperforms these baselines in handling motion blur and generating sharp novel views and novel poses. We visualize the out-of-distribution pose animation on ZJU-MoCap-Blur with pose sequences from AMASS [47] and AIST++ [37] in Fig. 5. This demonstrates our model’s generalization ability to extreme out-of-distribution poses. We use an off-the-shelf [27] pose estimator for the in-the-wild Real-Human-Blur dataset, yet we still achieve satisfying results. Refer to the supplementary for more results e.g. the qualitative comparisons on the baseline models with video deblurring preprocessing, and the effectiveness of the video deblurring method on our inputs.
4.4 Quantitative Results
@@
The quantitative results on ZJU-MoCap-Blur are reported in Sec. 4 and Sec. 4. The baseline model in Sec. 4 is trained on original motion-blurred inputs, and the inputs in Sec. 4 are pre-processed by video deblurring. Overall, our proposed approach performs better on PSNR and SSIM, and outperforms all the baselines on LPIPS, which is more informative in a monocular setting [62]. Video deblurring preprocessing improves the baselines only by a small margin, because of its focus on blur from camera motion and this 2D image-space paradigm lacks 3D consistency. Our method is capable of fast training and renders at a real-time rendering frame rate. Human motion trajectory modeling and pose-dependent fusion do not increase too much training time and do not add any rendering cost. Although Gauhuman [16] performs well on the original ZJU-MoCap dataset, discarding non-rigid deformation and implementing a new optimization pipeline for fast convergence result in its poor performance on the motion-blurred inputs. We ensure all methods are evaluated under the same version of SSIM and LPIPS calculators because different versions lead to numerical differences. GauHuman [16] is trained for 7k iterations instead of 3k as original paper to get the best performance.
4.5 Ablation Study
@@
Deblur-Avatar addresses the motion blur by predicting human pose sequences from human movements and incorporating the motion blur formation into 3DGS training by pose-dependent fusion. Therefore we conduct ablation studies on the human motion trajectory modeling module and the fusion module. As shown in Fig. 6 and Tab. 3, our framework works best when all the components are applied. We also conduct experiments on the number of virtual human poses in the exposure time in Fig. 5(b), the results show that increasing virtual poses is not always optimal if we aim to balance efficiency and quality.
@@
Table 3: Ablation study on the effectiveness of our proposed modules. We evaluate the six scenes of ZJU-MoCap-Blur, and we present the performance on average.
@@
Table 4: Ablation study on the number of virtual human poses n. We evaluate on two scenes. The results indicate that performance does not necessarily improve with increasing number n, but the increasing number will prolong the training time.
@@
5 Conclusion
We present Deblur-Avatar, a novel framework for sharp reconstruction of clothed human avatars from motion-blurred monocular video. To tackle motion blur from human movements in video capture, we model the human motion trajectories in 3DGS. Each timestamp is regarded as a sequence of sharp images captured within exposure time, and a sequence of human poses is predicted based on the input parameters. As human movements rarely involve the whole body, we predict masks to indicate the degradation regions within an image, enabling effective training on both blurred regions and sharp regions. Extensive experiments show that our real-time rendering method produces high-quality sharp avatars compared to state-of-the-art works.
Limitations. (1) We rely solely on a single input pose for trajectory modeling. A potential future direction is to explore inter-frame relationships to improve pose trajectory representation. (2) Deblur-Avatar does not reconstruct the accurate geometry of the avatar, and a potential direction is to extract smooth geometry from the 3DGS human avatar model by incorporating mesh or regularizing normal map. (3) Like previous works, Deblur-Avatar may produce poorly with high-frequency details e.g. complex local cloth, and the refinement for these areas is yet to be explored.

### v3

@@ IV-B Baseline Comparisons
The baselines include two types: 1) state-of-the-art human avatar reconstruction methods, e.g., NeRF-based methods HumanNeRF [42] and ARAH [10], 3DGS-based methods GauHuman [56], GoMAvatar [59], 3DGS-Avatar [20] and GART [16] and 2) an image-space baseline that uses pre-trained video deblurring [84] for preprocessing and then trains the human avatar baselines with the deblurred inputs. These baselines are compared under the monocular setup on ZJU-MoCap-Blur and Real-Human-Blur. All experiments are conducted on an NVIDIA RTX 3090 GPU.
IV-C Qualitative Results
@@
In Fig. 3, we present a comparative evaluation of our method against several state-of-the-art human avatar modeling approaches on the ZJU-MoCap-Blur dataset. As shown in the figure, current approaches struggle to reconstruct fine-grained details from motion-blurred inputs. In Fig. 4, we show the novel pose synthesis results on the Real-Human-Blur dataset. One can observe that current human avatar modeling methods cannot recover sharp details from motion blur, and our method outperforms these baselines in handling motion blur and generating sharp novel views and poses. We visualize the out-of-distribution pose animation on ZJU-MoCap-Blur with pose sequences from AMASS [99] and AIST++ [100] in Fig. 6. This demonstrates our model’s generalization ability to extreme out-of-distribution poses. We use an off-the-shelf [97] pose estimator for the in-the-wild Real-Human-Blur dataset, yet we still achieve satisfying results. We also present the qualitative comparisons on the baseline models with video deblurring preprocessing, as shown in Fig. 5.
IV-D Quantitative Results
@@
The quantitative results on ZJU-MoCap-Blur are reported in Section IV and Section IV. The baseline model in Section IV is trained on original motion-blurred inputs, and the inputs in Section IV are pre-processed by video deblurring. Overall, our proposed approach performs better on PSNR and SSIM, and outperforms all the baselines on LPIPS, which is more informative in a monocular setting [20]. Video deblurring preprocessing improves the baselines only by a small margin, because of its focus on blur from camera motion and this 2D image-space paradigm lacks 3D consistency. Our method is capable of fast training and renders at a real-time rendering frame rate. Human motion trajectory modeling and pose-dependent fusion do not increase too much training time and do not add any rendering cost. Although Gauhuman [56] performs well on the original ZJU-MoCap dataset, discarding non-rigid deformation and implementing a new optimization pipeline for fast convergence result in its poor performance on the motion-blurred inputs. We ensure all methods are evaluated under the same version of SSIM and LPIPS calculators because different versions lead to numerical differences. GauHuman [56] is trained for 7k iterations instead of 3k as original paper to get the best performance.
IV-E Ablation Study
@@
Our proposed method addresses the motion blur by predicting human pose sequences from human movements and incorporating the motion blur formation into 3DGS training by pose-dependent fusion. Therefore we conduct ablation studies on the human motion trajectory modeling module and the fusion module. As shown in Fig. 7 and Table III, our framework works best when all the components are applied. We also conduct experiments on the number of virtual human poses in the exposure time in Table IVb, the results show that increasing virtual poses is not always optimal if we aim to balance efficiency and quality.
The results of our ablation study on trajectory representations are presented in Table Vb. We conduct on two settings: (1) optimizing $\theta_{\text{start}}$ and $\theta_{\text{end}}$ and interpolate the human poses by linear interpolation, and (2) utilizing a high-order cubic B-spline that jointly optimizes four control knots to depict human motions. The results demonstrate that Spherical Linear Interpolation adequately represents the human motion trajectory.
We also explore whether adding a learnable interpolation will further improve the performance in Table VIb. The results show that the current interpolation method is adequate for human motion trajectory modeling.
@@
Table III: Ablation study on the effectiveness of our proposed modules. We evaluate the six scenes of ZJU-MoCap-Blur, and we present the performance on average.
@@
Table IV: Ablation study on the number of virtual human poses n. We evaluate on two scenes. The results indicate that performance does not necessarily improve with increasing number n, but the increasing number will prolong the training time.
@@
Table V: Ablation study on the trajectories representations. We evaluate four scenes from ZJU-MoCap-Blur. The best performance of each scene is boldfaced on each row.
[…]
377 30.36 0.9765 26.69
386 33.75 0.9770 29.37
387 28.61 0.9629 40.39
392 30.45 0.9697 38.49
[…]
377 30.19 0.9758 27.58
386 33.86 0.9772 29.49
387 28.66 0.9630 40.62
392 30.43 0.9694 39.33
Table VI: Ablation study on the learnable trajectories representations. We evaluate two scenes from ZJU-MoCap-Blur. The best performance of each scene is boldfaced on each row.
[…]
377 30.36 0.9765 26.69
386 33.75 0.9770 29.37
[…]
377 30.31 0.9759 27.81
386 33.85 0.9773 29.51
IV-F Computational Efficiency
In terms of resource requirements, we conduct experiments to calculate the training and inference costs. Our method costs more memory than 3DGS baselines due to the increased complexity of modeling motion blur, but saves more memory than NeRF baselines in Table VII. This trade-off highlights the balance our method strikes between performance and computational cost, offering a practical alternative for high-quality human avatar modeling under motion blur.
Table VII: Training and inference efficiency.
Method HumanNeRF Arah GauHuman GART 3DGS-Avatar GoMA Ours
Training Memory (G) 11.3 11.9 2.0 2.5 6.4 4.4 7.3
Inference Memory (G) 18.7 19.2 0.9 3.3 3.1 3.2 2.7
IV-G Difference With Current Deblurring Methods
As shown in Section IV and Fig. 5, the current state-of-the-art video deblurring method [84] only improves the human avatar modeling baselines by a small margin. Therefore, we conduct qualitative experiments solely on the performance of the video deblurring method. As shown in Fig. 8, the current deblurring pipeline works well on motion blur due to camera movements. However, the pipeline fails to handle motion blur from human movements, because the blur pattern of the input is different.
Figure 8: Pre-deblurring results of state-of-the-art video deblurring method [84]. On the first row we show the data where this method performs well. Current video deblurring paradigms are fitting for motion blur from camera shake. However, the motion blur in the human avatar modeling task is primarily from human movements, and current methods perform poorly.
V Conclusion
We present a novel framework for sharp reconstruction of clothed human avatars from motion-blurred monocular video. To tackle motion blur from human movements in video capture, we model the human motion trajectories in 3DGS. Each timestamp is regarded as a sequence of sharp images captured within exposure time, and a sequence of human poses is predicted based on the input parameters. As human movements rarely involve the whole body, we predict masks to indicate the degradation regions within an image, enabling effective training on both blurred regions and sharp regions. Extensive experiments show that our real-time rendering method produces high-quality sharp avatars compared to state-of-the-art works.
Limitations. (1) We rely solely on a single input pose for trajectory modeling. A potential future direction is to explore inter-frame relationships to improve pose trajectory representation. (2) The proposed method does not reconstruct the accurate geometry of the avatar, and a potential direction is to extract smooth geometry from the 3DGS human avatar model by incorporating mesh or regularizing normal map. (3) Like previous works, our method may produce poorly with high-frequency details, e.g., complex local cloth, and the refinement for these areas is yet to be explored.